### v3

@@ Abstract
Question answering (QA) is an actively studied topic, being a core natural language processing (NLP) task that needs to be addressed before achieving Artificial General Intelligence (AGI). However, the lack of QA datasets in specific domains and languages hinders the development of robust AI models able to generalize across various domains and languages. To this end, we introduce MedQARo, the first large-scale medical QA benchmark in Romanian, alongside a comprehensive evaluation of state-of-the-art (SOTA) large language models (LLMs). We construct a high-quality and large-scale dataset comprising 105,880 QA pairs related to cancer patients from two medical centers. The questions regard medical case summaries of 1,242 patients, requiring either keyword extraction or reasoning to be answered correctly. MedQARo is the result of a time-consuming manual annotation process carried out by seven physicians specialized in oncology or radiotherapy, who spent a total of about 3,000 work hours to generate the QA pairs. Our benchmark contains both in-domain and cross-domain (cross-center and cross-cancer) test collections, enabling a precise assessment of generalization capabilities. We experiment with four open-source LLMs from distinct families of models on MedQARo. Each model is employed in two scenarios, namely one based on zero-shot prompting and one based on supervised fine-tuning. We also evaluate two state-of-the-art LLMs exposed only through APIs, namely GPT-5.2 and Gemini 3 Flash. Our results show that fine-tuned models significantly outperform zero-shot models, clearly indicating that pretrained models fail to generalize on MedQARo. Our findings demonstrate the importance of both domain-specific and language-specific fine-tuning for reliable clinical QA in Romanian. We publicly release our dataset and code at https://github.com/ana-rogoz/MedQARo.
@@ 1 Introduction
The development of robust question answering (QA) systems is an important goal in natural language processing (NLP), representing one of the prerequisites for building systems that reach human-level intelligence. Despite the significant advances in the area of Large Language Models (LLMs) [OpenBioLLMs, touvron2023llama, jiang2023mistral7b, alpaca, kopf2024openassistant], which are able to perform remarkably well across various QA tasks [brown2020language, chowdhery2023palm, rajpurkar-etal-2016-squad, hendrycksmeasuring, openai2023gpt4, singhal2023large, singhal2023towards], QA remains a challenging task within specialized domains, such as the medical field, or in low-resource languages, such as Romanian. For instance, the task of medical QA comes with unique challenges due to the complexity of clinical narratives and the necessity for deep understanding of medical terms. Moreover, there is a soaring demand for high accuracy, as mistakes can have serious consequences in daily medical practice assisted by AI agents. While large-scale datasets such as PubMedQA [jin-etal-2019-pubmedqa], emrQA [pampari-etal-2018-emrqa], and MedQA [jin2020diseasedoespatienthave] have driven significant progress in English medical QA, the availability of similar resources in other languages, especially underrepresented ones, remains limited. This hinders the development of LLMs that are robust to language and domain shift. To the best of our knowledge, for the Romanian language, there are no prior efforts in the realm of large-scale QA resources for the medical domain.
Figure 1: An image illustrating our dataset creation and model benchmarking stages.
To address this gap, we introduce MedQARo, the first large-scale benchmark for Medical Question Answering in Romanian. As illustrated in Figure 1, the benchmark creation involves two phases, (i) manual data collection and annotation, and (ii) model benchmarking. The dataset comprises 105,880 high-quality QA pairs from real-world clinical records of 1,242 oncology patients (796 patients with breast cancer, 215 patients with lung cancer, and 231 patients with other cancers). The QA pairs are the results of a manual annotation process carried out by physicians specialized in oncology and radiotherapy, who collectively spent nearly 3,000 hours to generate the pairs, ensuring both linguistic quality and medical correctness. MedQARo includes 76,416 QA pairs about breast cancer patients, 26,230 about lung cancer patients and 3,234 about patients with other cancer types, with questions grounded in medical case summaries (epicrises).
To prevent data leakage across splits and support fair model development and evaluation, we provide official training, validation, and test splits at the patient level, i.e. a medical record can appear in only one of the splits. Moreover, we provide both in-domain and cross-domain test collections, where the cross-domain test set comprises patients from a distinct medical center, who suffer from five cancer types (head and neck, melanoma, renal cell carcinoma, bladder cancer, hepatocellular carcinoma) that are not available during training. Hence, MedQARo enables a precise assessment of generalization capabilities across different medical centers and cancer types. Oncologists, especially residents, could greatly benefit from using LLMs to extract and reason about the information included in epicrises, when they need to recommend diagnostic tools or treatment lines for their patients. MedQARo provides the necessary means to develop such LLMs, that could be deployed in daily medical practice. We benchmark four models on MedQARo, namely two LLMs specialized on Romanian (RoLLaMA2-7B and RoMistral-7B) [masala2024vorbecstiromanecsterecipetrain], a long-context LLM (Phi-4-mini-instruct) [phi4mini], and an LLM tuned on biomedical data (LLaMA3-OpenBioLLM-8B) [OpenBioLLMs]. Each model is assessed in both zero-shot and supervised fine-tuning setups, while testing various prompt formats. In addition, we assess two state-of-the-art LLMs in the zero-shot setting, which are only accessible through paid APIs, namely GPT-5.2 and Gemini 3 Flash. Our results show that all LLMs significantly benefit from fine-tuning, whereas zero-shot variants perform poorly. These findings emphasize the critical importance of both domain and language adaptation for reliable clinical QA in Romanian. Among all evaluated models, the fine-tuned Phi-4-mini-instruct achieves the best performance, significantly outperforming other approaches. However, the best-performing model achieves an F1 score of only $0.667$ on the in-domain test set, suggesting that MedQARo is a challenging dataset, highlighting the need for the development of more robust models, capable of better generalizing to specialized domains and low-resource languages.
In summary, our contributions are fivefold:
  • We release the dataset and code to reproduce the results in a public repository, where patient sensitive data is fully anonymized.
2 Related Work
@@
English and multi-lingual QA datasets. Question answering emerged as a fundamental task in NLP, with several datasets being developed to support research across various domains and languages. Popular QA datasets focusing on the English language, such as SQuAD [rajpurkar-etal-2016-squad] and Natural Questions [kwiatkowski-etal-2019-natural], established the foundation for extractive and generative QA systems. These datasets showed the importance of large-scale high-quality annotations for training robust QA models. Later on, given the need for cross-language understanding, more and more datasets covering several languages started to arise. Some notable multilingual efforts include XQuAD [artetxe-etal-2020-cross], which extends SQuAD to 11 languages, and MLQA [lewis-etal-2020-mlqa], which provides QA datasets in seven languages with cross-lingual evaluation capabilities. Furthermore, TyDi QA [clark-etal-2020-tydi] advanced multilingual QA by covering structurally different languages and focusing on information-seeking questions. However, these efforts did not include Romanian among the target languages.
@@
Medical QA datasets. Medical QA represents a challenging domain that requires deep understanding of a broad range of aspects, such as clinical terminology and medical reasoning. Several English medical QA datasets have been already developed to tackle the particularities of the medical domain. For instance, MedQA [jin2020diseasedoespatienthave] provides over 12,000 multiple-choice questions from medical licensing exams across three countries, testing complex clinical reasoning abilities. Additionally, the emrQA dataset [pampari-etal-2018-emrqa] introduces one of the largest QA medical resources with over 455,000 question-answer pairs automatically generated from electronic medical records. PubMedQA [jin-etal-2019-pubmedqa] focuses on biomedical literature comprehension, with questions derived from research article titles and abstracts. MedMCQA [pmlr-v174-pal22a] expands medical QA to include Indian medical entrance exams, with 193,100 questions across 21 medical subjects, demonstrating the importance of diverse medical knowledge sources. Despite its data source (Indian medical exams), the QA pairs contained by MedMCQA are in English.
In contrast to the significant advances in English medical QA, the landscape of non-English medical QA datasets remains scarce. To the best of our knowledge, we are the first to introduce a large-scale QA dataset specifically tailored for the medical domain in Romanian. MedQARo contains original clinically-grounded content rather than translations, addressing a critical gap in current literature.
@@
Romanian QA datasets. For the Romanian language, a few QA datasets have been independently developed. Two such datasets focus on domain specific understanding. One is the JuRo dataset [craciun-etal-2025-graf], which provides a collection of Romanian juridical texts with associated questions, and the other is the RoITD dataset [nicolae2021roitd], which addresses technical domain questions in the IT sector. Additionally, the LiRo benchmark [liro2021], which provides a comprehensive evaluation platform for Romanian NLP tasks, contributed with a subset of XQuAD [artetxe-etal-2020-cross] for Romanian, consisting of translated versions of English questions. For advanced biology, dima-etal-2024-roqllama published a small-scale dataset comprising 4,127 single-choice questions. Existing datasets for Romanian QA are typically small (containing between 1,000 and 11,000 QA pairs). With 105,880 QA pairs, MedQARo is one order of magnitude larger than the largest dataset for Romanian QA.

### v4

@@ Abstract
We introduce MedQARo, the first large-scale medical QA benchmark in Romanian, alongside a comprehensive evaluation of state-of-the-art large language models (LLMs). We construct a high-quality and large-scale dataset comprising 105,880 QA pairs about cancer patients from two medical centers. The questions regard medical case summaries of 1,242 patients, requiring both keyword extraction and reasoning. Our benchmark contains both in-domain and cross-domain (cross-center and cross-cancer) test collections, enabling a precise assessment of generalization capabilities. We experiment with four open-source LLMs from distinct families of models on MedQARo. Each model is employed in two scenarios: zero-shot prompting and supervised fine-tuning. We also evaluate two state-of-the-art LLMs exposed only through APIs, namely GPT-5.2 and Gemini 3 Flash. Our results show that fine-tuned models significantly outperform zero-shot models, indicating that pretrained models fail to generalize on MedQARo. Our findings demonstrate the importance of both domain-specific and language-specific fine-tuning for reliable clinical QA in Romanian.
@@ Introduction
The development of robust question answering (QA) systems is an important goal in natural language processing (NLP), representing one of the prerequisites for building systems that reach human-level intelligence. Despite the significant advances in the area of Large Language Models (LLMs) [OpenBioLLMs, touvron2023llama, jiang2023mistral7b, alpaca, kopf2024openassistant], which are able to perform remarkably well across various QA tasks [brown2020language, chowdhery2023palm, rajpurkar-etal-2016-squad, hendrycksmeasuring, openai2023gpt4, singhal2023large, singhal2023towards], QA remains a challenging task within specialized domains, such as the medical field, or in low-resource languages, such as Romanian. For instance, the task of medical QA comes with unique challenges due to the complexity of clinical narratives and the necessity for deep understanding of medical terms. Moreover, there is a soaring demand for high accuracy, as mistakes can have serious consequences in daily medical practice assisted by AI agents.
@@
Question answering emerged as a fundamental task in NLP, with several datasets being developed to support research across various domains and languages. Popular QA datasets focusing on the English language, such as SQuAD [rajpurkar-etal-2016-squad] and Natural Questions [kwiatkowski-etal-2019-natural], established the foundation for extractive and generative QA systems. These datasets showed the importance of large-scale high-quality annotations for training robust QA models. Later on, given the need for cross-language understanding, more and more datasets covering several languages started to arise. Some notable multilingual efforts include XQuAD [artetxe-etal-2020-cross], which extends SQuAD to 11 languages, and MLQA [lewis-etal-2020-mlqa], which provides QA datasets in seven languages with cross-lingual evaluation capabilities. Furthermore, TyDi QA [clark-etal-2020-tydi] advanced multilingual QA by covering structurally different languages and focusing on information-seeking questions. However, these efforts did not include Romanian among the target languages.
@@
Medical QA represents a challenging domain that requires deep understanding of a broad range of aspects, such as clinical terminology and medical reasoning. Several English medical QA datasets have been already developed to tackle the particularities of the medical domain. For instance, MedQA [jin2020diseasedoespatienthave] provides over 12,000 multiple-choice questions from medical licensing exams across three countries, testing complex clinical reasoning abilities. Additionally, the emrQA dataset [pampari-etal-2018-emrqa] introduces one of the largest QA medical resources with over 455,000 question-answer pairs automatically generated from electronic medical records. PubMedQA [jin-etal-2019-pubmedqa] focuses on biomedical literature comprehension, with questions derived from research article titles and abstracts. MedMCQA [pmlr-v174-pal22a] expands medical QA to include Indian medical entrance exams, with 193,100 questions across 21 medical subjects, demonstrating the importance of diverse medical knowledge sources. Despite its data source (Indian medical exams), the QA pairs contained by MedMCQA are in English.
While large-scale datasets such as PubMedQA [jin-etal-2019-pubmedqa], emrQA [pampari-etal-2018-emrqa], and MedQA [jin2020diseasedoespatienthave] have driven significant progress in English medical QA, the availability of similar resources in other languages, especially underrepresented ones, remains limited. This hinders the development of LLMs that are robust to language and domain shift. To address this gap, we introduce MedQARo, the first large-scale benchmark for Medical Question Answering in Romanian. MedQARo contains original clinically-grounded content rather than translations, addressing a critical gap in current literature.
@@
For the Romanian language, a few QA datasets have been independently developed. Two such datasets focus on domain specific understanding. One is the JuRo dataset [craciun-etal-2025-graf], which provides a collection of Romanian juridical texts with associated questions, and the other is the RoITD dataset [nicolae2021roitd], which addresses technical domain questions in the IT sector. Additionally, the LiRo benchmark [liro2021], which provides a comprehensive evaluation platform for Romanian NLP tasks, contributed with a subset of XQuAD [artetxe-etal-2020-cross] for Romanian, consisting of translated versions of English questions. For advanced biology, dima-etal-2024-roqllama published a small-scale dataset comprising 4,127 single-choice questions. Existing datasets for Romanian QA are typically small (containing between 1,000 and 11,000 QA pairs). With 105,880 QA pairs, MedQARo is one order of magnitude larger than the largest dataset for Romanian QA.